OpenStudy (anonymous):

Find the slope of the line passing through the points (2, 5) and (0, 4).

OpenStudy (anonymous):

use the formula.

OpenStudy (lalaly):

\[\underline{slope}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

k. point slope form here! 5 - 4 = m(2-0) now simplify 1 = m(2) now divide by 2 1/2 = m now you know that your slope is 1/2 :)

OpenStudy (sasogeek):

slope=(5-4)/(2-0)=1/2